OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 2x - 9y = 1 x - 4y = 1 (1, 5) (-5, -1) (5, 1) (-5, 1)

OpenStudy (anonymous):

solution is (5, 1)

OpenStudy (anonymous):

Solve the second equation for x: x=4y+1 (*) Sub (*) into first equatiuon to get a equation in y 2(4y+1)-9y=1 Solve this one for y 8y+2-9y=1 -y=-1 y=1 (**) Sub (**) into (*) to get x x=4(5)+1=5 Therefor solution is as raheen gave

OpenStudy (amistre64):

substitution is basic enough; i like cramers for practice tho :) 2x - 9y = 1 x - 4y = 1 y = 2-1 / -8+9 = 1/1 = 1 x = -9+4 / 8-9 = -5 / -1 = 5

OpenStudy (anonymous):

Solve the second equation for x: x=4y+1 (*) Sub (*) into first equatiuon to get a equation in y 2(4y+1)-9y=1 Solve this one for y 8y+2-9y=1 -y=-1 y=1 (**) Sub (**) into (*) to get x x=4(5)+1=5 Therefor solution is as raheen gave

OpenStudy (anonymous):

Definitley (5,1)